OpenStudy (he66666):

Differential equations: y' = y(3 − ty) Describe how solutions appear to behave as t increases and how their behavior depends on the initial value y0 ("0" is a subscript) when t = 0. How do you solve this question? I don't think the equation is linear so I am confused as to how to solve it.. Can someone help me please?

OpenStudy (lgbasallote):

i think it's more of a bernoulli

OpenStudy (anonymous):

it is a bernoulli...make it linear by dividing the differential equation by y^2...and put 1/y = z and then it becomes linear in z..

OpenStudy (anonymous):

I hope you got it..proceed as a linear differential therefurther...

OpenStudy (he66666):

Oh I see, but how do you divide y' by y^2?

OpenStudy (anonymous):

when you put z=1/y, you get by differentiating dz/dt = -1/y^2 x dy/dx So, -dz/dt = (dy/dx)(1/y^2)

OpenStudy (he66666):

I got (1/y^2)(dy/dt) = (1/y)[(3/y^2)-(t/y)] After I put 1/y as z, (z^2)(-1/z^2)(dy/dt) = z +3z -tz -(dy/dt) = z-tz+3z^2

OpenStudy (he66666):

Is this right?

OpenStudy (anonymous):

just simply write DE as \[\Large y'=3y-ty^2\] now multiply whole equation with \[\Large y^{-2}\] \[\Large y^{-2}y'=3y^{-1}-t\] let \[\Large v=y^{-1}\] \[\Large v'=-y^{-2}y'\] or \[\Large -v'=y^{-2}y'\] just substitute in above for ^-2y' and y^-1 \[\Large -v'=2v-t\] or \[\Large v'+3v=t\] this is now Linear equation just find the integrating factor and solve it .

OpenStudy (anonymous):

does that help you ?

OpenStudy (anonymous):

there is typo mistake from third last line it is \[\Large -v=3v-t\]

OpenStudy (he66666):

Thanks a lot!